OpenStudy (fibonaccichick666):

Inverse Laplace transform question: Find Y \[L(y)=\frac{s-1}{(s^2-1)((s-1)^2+1)}\] I'm hitting a road block here

OpenStudy (fibonaccichick666):

@zepdrix

OpenStudy (fibonaccichick666):

wolfram is telling me it should be immaginary

OpenStudy (fibonaccichick666):

@Kainui, @ganeshie8 not sure if this is your territory

hartnn (hartnn):

have you considered frequency shifting property ? \(e^{at}f(t) <-> F(s-a)\)

OpenStudy (fibonaccichick666):

I'm not sure how that would work here

OpenStudy (fibonaccichick666):

(I've never heard that particualr phrase either)

hartnn (hartnn):

\(\Large L^{-1}[\dfrac{s-1}{(s^2-1)((s-1)^2+1)}] = e^t L^{-1}[\dfrac{s}{s(s+2)((s)^2+1)}]\)

hartnn (hartnn):

after that normal partial fraction would work :)

OpenStudy (fibonaccichick666):

ahh ok Thanks, I think that will work! I'll check back once I give it a shot

hartnn (hartnn):

\(\Large \dfrac{A}{s+2} + \dfrac{Bs+C}{s^2+1}\)

hartnn (hartnn):

sure, ask if any more doubts, now or later :)

OpenStudy (fibonaccichick666):

Thank you! that was it! Quick question if you shift it a times would it be \(e^{at}\)?

hartnn (hartnn):

thats correct. if you see s-a in the s domain, you can plug out e^(at) in t domain

OpenStudy (fibonaccichick666):

thank you!!!!! I totally forgot about this